# Improving a Proportional Integral Controller with Reinforcement Learning on a Throttle Valve Benchmark

### v1

@@ I Introduction
Throttle valves are essential components in various industrial processes such as chemical plants, oil refineries, and power generation. Accurate control of these valves is crucial for maintaining the optimal flow rate of fluids and thus ensuring the efficient functioning of the entire system. However, controlling throttle valves is complex due to their non-linear behavior, stochasticity, and asymmetric hysteresis. These factors make it challenging to design an optimal controller that can handle the complexities of the valve system.
In this work, we focus on the regulation of a butterfly valve for car engines, crucial for modulating the amount of air supplied to the combustion chamber by adjusting a disc’s rotation angle. Given the complexities induced by static friction and the non-linearities of the dynamics, the control community has explored various advanced strategies. For example, non-linear control approaches include an adaptive pulse control method leveraging a non-linear dynamic model accounting for friction and aerodynamic torque [1], a discrete-time sliding mode control for robust tracking [2], and a non-linear approach combining a Proportional-Integral-Derivative (PID) controller with a feedback compensator [3]. Additionally, adaptive control techniques [4, 5, 6], Linear Parameter-Varying (LPV) modeling, and mixed constrained $H_{2}/H_{\infty}$ control strategies [7] have been proposed. Several works also exploit deep learning techniques. For instance, [8] presents a neural network-based sliding mode controller for electronic throttle valves. A control strategy using RL algorithms is advocated in [9] where the valve is controlled using the Deep Deterministic Policy Gradient (DDPG) [10] algorithm. However, the application of the algorithm is limited to simulations.
@@
Only a few recent works combine Optimal Control (OC) and RL to reduce the number of data required by RL agents to learn a good policy [18].
The approaches of [19] and [20] propose a switching mechanism between an LQR controller and an RL agent, testing these approaches in simulated environments such as a pendulum and Acrobot. Some approaches embed a feedback controller within an RL policy to enhance sample efficiency [21, 22]. However, these methods do so without restraining the search space for the control input, nor without modifying the learning of the $Q$-function, which may be essential to avoid extrapolation errors [23, 24]. Similar to these previous methods, we aim to combine the strengths of OC and RL to build a controller for the throttle valve. For this, we adapt the recent area of Reinforcement Learning with Guides [25, 26, 27] to our setting, resulting in an algorithm that optimizes a Proportional Integral (PI) controller to guide the RL agent in navigating the control space efficiently. By reducing the search space for learning, the PI guide minimizes the data requirements and accelerates the learning process. We stress that our method is fundamentally different than those proposed by [28, 29, 30], which use RL to tune the gains of a feedback PI controller. Here, the PI controller is fixed and is used to guide the RL agent to improve its sample efficiency.
@@ II-B A non-linear stochastic system
This type of valve is particularly known for its non-linear stochastic dynamics with asymmetric hysteresis, which we confirm in the following experiment. We select an input range from $0\%$ to $\text{PWM}_{\text{MAX}}\%$ and generate an increasing then decreasing sequence of steps covering this range, by increments of $5\%$, and apply it to the valve. Despite the valves coming from the same commercial reference, we notice that the required value $\text{PWM}_{\text{MAX}}$ to turn the valve’s angular plate to 00 degrees is different from one valve to another [6]. It is $\text{PWM}_{\text{MAX}}=0.65\%$ for Valve $1$, $\text{PWM}_{\text{MAX}}=0.70\%$ for Valve $2$ and $\text{PWM}_{\text{MAX}}=0.45\%$ for Valve $3$. The experiment is repeated $5$ times, and we present the resulting data in Figure 2. It can be seen that for all valves we have an asymmetric hysteresis as well as the stochastic nature of the dynamics. For a given PWM input, the resulting angles vary significantly between each trial.
@@ II-C Problem statement and notations
Considering the experimental set-up and the stochastic nature of the valves’ dynamics, the problem is modeled as a discounted Markov Decision Process (MDP) $\left(\mathcal{X},\mathcal{U},c,P,\gamma\right)$. Let $\Delta(\cdot)$ be the simplex on any given space ($\cdot$), and define $c:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ the cost function, $P:\mathcal{X}\times\mathcal{U}\to\Delta(\mathcal{X})$ the transition probabilities, and $\gamma\in\left[0,1\right)$ the discount factor. At each time step $t$, the agent receives an observation $x_{t}\in\mathcal{X}$ and selects a control $u_{t}\in\mathcal{U}$. The system is then updated with $x_{t+1}\sim P\left(\cdot|x_{t},u_{t}\right)$.
@@
The goal of this work is to develop a throttle valve controller $\pi:\mathcal{X}\to\Delta(\mathcal{U})$ that sets the angle of the valve’s plate to a chosen angle from any initial position. The controller must not only perform optimally in nominal operating conditions by minimizing cumulative costs but also demonstrate robustness and stability under various real-world scenarios. Specifically, the controller must be able to reject external disturbances that may affect both the input and output signals, while exhibiting minimal overshoot. Bearing this information in mind, we associate the different components of the MDP with the valve. At time $t$, let $\alpha_{t}$ be the angle of the valve’s plate and $\alpha_{\text{ref}_{t}}$ be the (desired) reference angle.
Furthermore:
@@
  • the controller must be optimal in all states and take into account long-term performance, so the discount factor $\gamma$ is set to the high value of $0.99$;
@@ III-B Approximate Policy Iteration
This cycle of evaluation and improvement is commonly solved approximately with gradient descent and continues until the policy converges to an optimal one. This framework, not requiring prior knowledge from the environment beyond interaction data, can be applied as such to any system as long as it possesses the Markov property. Nonetheless, the absence of initial information makes it difficult to collect meaningful data at the early stages of learning. This leads to poor estimates of the $Q$ function and the policy. As a result, agents necessitate extensive interactions to build accurate estimates and derive a near-optimal policy.
@@ IV Combining both methods
Figure 3: Comparison of the different agents under different scenarios when all agents have access to the real outputs. Both figures display the evolution of the angles over time with a reference change every $X$ seconds, set to $5$ seconds for the left figures and to $2.5$ seconds for the right figures.
@@
Unlike the traditional RL policy $\pi_{\text{RL}}^{\theta}$ that provides a control input $u$, the perturbation $\xi^{\phi}_{\text{RL}}$ aims solely to enhance the PI controller’s performance. Thus, rather than defining the perturbation $\xi^{\phi}_{\text{RL}}$ over the entire control space $\mathcal{U}$, it is constrained to a smaller subset $S(\mathcal{U})$. This confines the search for optimal control to a vicinity around the PI controller, where the optimal control is more likely to be found, thereby improving the algorithm’s sample efficiency. In practice, to reduce the search space that is originally $[0,0.6]$ or $[0,0.8]$ and keeping the differentiability property of the policy, we introduce a multiplicative factor $\eta\in[0,1]$, which scales the control input by a factor $\eta$.
The perturbation is learned similarly to a control input in traditional RL, that is by minimizing its associated $Q^{\omega}$ function. We stress here that the APE step has to be modified to take into account the policy change. Note that the expectation in the Bellman operator is with respect to $\pi^{\phi_{k}}_{\text{PI-RL}}$, not $\pi_{\theta_{k}}$. This avoids the overestimating problem that may be encountered when the $Q$-function estimates observation-control pairs that are not described by the data set $\mathcal{D}$ [32, 27]. The process becomes the PI-Approximate Policy Iteration, where the APE-API steps are modified to be PI-APE and PI-API, defined as:
@@ V Comparison of the different controllers on the experimental test benches
Figure 4: Comparison of the different agents under noisy outputs or noisy controls with different standard deviations. Each bar represents the Mean-Squared-Error (MSE) between the reference and the angle of the valve.
We conduct a series of experiments on various setups. We first perform a nominal comparison between the controllers, where we test the controllers’ performance on a simple task: tracking a reference signal with sufficient time to adjust. We then test them on more challenging tasks, including adapting to quick reference changes and robustness with respect to noise in the control input or in the output observation.
@@ V-A Nominal comparison and reaction to quick reference changes
In all scenarios, we observe that the PI controller, while efficient, encounters difficulties in accurately tracking the reference signal. The controller requires time to accumulate errors before adjusting to the reference. Due to the system’s complexity and the dry friction effect, it may cause the valve’s plate to overshoot the target. The PI controller would then need additional time to adjust, and potentially select overly aggressive inputs once more. This results in oscillations around the reference angle, particularly evident in the first valve that seems the most complicated to control. The other valves also present such oscillations, especially for low-angle references. This reveals the need for more sophisticated controllers and not solely relying on a linear policy.
@@
On the other hand, both RL-based agents demonstrate remarkable success in precisely tracking the reference signal in all scenarios, including the most challenging ones. For example, they can track all extreme reference values with a change every 2.5 seconds, which is twice as fast as the rate they were trained for. This level of performance is also observed for low-angle references that proved to be challenging for the PI controller. The controls provided by the RL-based agents are stable in all settings, with no instances of unstable behavior observed. In some cases, a slight tracking bias is observed where the agents reach a point close to but not precisely at the reference. This can be attributed to the RL agents’ goal of minimizing the $Q$ function, which is an expectation over the system’s stochastic dynamics. To address this, a more appropriate RL formulation may be needed, which considers higher-order moments of the cumulative future cost distribution or steady-state errors in a refined way. We leave this for future work. Despite these minor biases, both RL-based agents demonstrate a near-optimal performance in all tasks for all valves.
@@ V-B Robustness to external noise
Figure 5: Learning curves of the different RL-based agents for the $3$ valves. We train both agents during $2500$ episodes, representing $3$ hours with our computer without the need to use a GPU.
@@
In comparison to the PI controller, both RL-based agents demonstrate superior performance over the PI controller in all scenarios. Once again, it is evident that the first valve is the most complicated to control as attested by the PI performances. It is also essential to note that, under specific conditions of noise, e.g. in Valve $2$ with a control noise with a standard deviation of $0.2$ and in Valve $1$ with an output noise of standard deviation $0.2$, the PI-RL agent exhibits a higher loss than the traditional RL agent. This result may be attributed to the fact that the PI-RL agent is centered around the PI controller which is inadequate in these scenarios. Consequently, the reduced search space is not sufficient to yield an optimal controller, and a more sophisticated approach may be needed to address the challenges posed by these specific noise conditions. Nevertheless, we note that the perturbation $\xi_{RL}^{\phi}$ can improve the PI controller up to a certain point.
@@ V-C Sample efficiency
As previously shown, both the PI-RL and classical RL agents achieve similar performances in controlling the throttle valve. One important advantage of the combined PI and RL approach over the classical RL approach is that it requires fewer data to learn even though the PI controller may not be the optimal one. This can be seen in Figure 5, where we plot the cumulative costs over the training process for both controllers for the $3$ different valves.
We observe that the PI-RL approach achieves a low cumulative cost faster than the RL approach, indicating that it was able to learn the control task more efficiently. This is because PI-RL leverages the prior knowledge of the PI controller to reduce the search space and guide the RL exploration, which helps to reduce the exploration time and improve the sample efficiency. After a sufficient amount of data, the traditional agent finally catches the performance of the combined agent. This is less obvious in Valve $3$, which seems to be the simplest valve to control as observed in Figure 3 and Figure 4. Indeed, the RL agent converges almost as fast as the PI-RL agent to the same near-optimal policy.
@@ VI Conclusion
In this work, we show that the combination of reinforcement learning and optimal control can be a powerful approach to designing near-optimal controllers for throttle valve systems. By adapting the recent area of Reinforcement Learning with Guides to the throttle valve system, we can build a near-optimal controller that reduces the data requirement of traditional RL agents and overcomes the limitations of the PI controller. Our approach is validated on three different valves with slightly different dynamics, all demonstrating the advantages and drawbacks of all controllers.
This provides evidence that the integration of OC techniques with RL can lead to significant improvements in data efficiency, making it a promising avenue for future research in other complex systems.

### v2

@@ I Introduction
Throttle valves are essential components in various industrial processes such as chemical plants, oil refineries, and power generation. Accurate control of these valves is crucial for maintaining the optimal flow rate of fluids and thus ensuring the efficient functioning of the entire system. However, controlling throttle valves is complex due to their non-linear behavior, stochasticity, and asymmetric hysteresis. These factors make it challenging to design a controller that can handle the complexities of the valve system.
In this work, we focus on regulating a butterfly valve for car engines, which modulates the air supply to the combustion chamber by adjusting the disc’s rotation angle. Given the complexities induced by static friction and the non-linearities of the dynamics, the control community has explored various advanced strategies. For example, non-linear control approaches include an adaptive pulse control method leveraging a non-linear dynamic model accounting for friction and aerodynamic torque [1], a discrete-time sliding mode control for robust tracking [2], and a non-linear approach combining a Proportional-Integral-Derivative (PID) controller with a feedback compensator [3]. Additionally, adaptive control techniques [4, 5, 6], Linear Parameter-Varying (LPV) modeling, and mixed constrained $H_{2}/H_{\infty}$ control strategies [7] have been proposed. Several works also exploit deep learning techniques. For instance, [8] presents a neural network-based sliding mode controller for electronic throttle valves. A control strategy using RL algorithms is advocated in [9] where the valve is controlled using the Deep Deterministic Policy Gradient (DDPG) [10] algorithm. However, the application of the algorithm is limited to simulations.
@@
Only a few recent works combine Control Theory (CT) and RL to reduce the number of data required by RL agents to learn a good policy [18].
The approaches of [19] and [20] propose a switching mechanism between an LQR controller and an RL agent, testing these approaches in simulated environments. Some approaches embed a feedback controller within an RL policy to enhance sample efficiency [21, 22]. However, these methods do so without restraining the search space for the control input, nor without modifying the learning of the $Q$-function, which may be essential to avoid extrapolation errors [23]. Similar to these previous methods, we aim to combine the strengths of CT and RL to build a controller for the throttle valve. For this, we adapt the recent area of Reinforcement Learning with Guides [24, 25, 26] to our setting, resulting in an algorithm that leverages a Proportional Integral (PI) controller to guide the RL agent. By reducing the search space for learning, the PI guide minimizes the data requirements and accelerates the learning process. We stress that our method is fundamentally different than those proposed by [27, 28, 29], which use RL to tune the gains of a feedback PI controller. Here, the PI controller is fixed and is used to guide the RL agent to improve its sample efficiency.
@@ II-B A non-linear stochastic system
This type of valve is particularly known for its non-linear stochastic dynamics with asymmetric hysteresis, which we confirm in the following experiment. We select an input range from $0\%$ to $\text{PWM}_{\text{MAX}}\%$ and generate an increasing then decreasing sequence of steps covering this range, by increments of $5\%$, and apply it to the valve. Despite the valves coming from the same commercial reference, we notice that the required value $\text{PWM}_{\text{MAX}}$ to turn the valve’s angular plate to 00 degrees is different from one valve to another [6]. It is $\text{PWM}_{\text{MAX}}=0.65\%$ for Valve $1$, $\text{PWM}_{\text{MAX}}=0.70\%$ for Valve $2$ and $\text{PWM}_{\text{MAX}}=0.45\%$ for Valve $3$. The experiment is repeated $5$ times, and we present the resulting data in Figure 2. An asymmetric hysteresis is observed for all valves, along with the stochastic nature of their dynamics. For a given PWM input, the resulting angles vary significantly between each trial.
@@ II-C Problem statement and notations
Considering the stochastic nature of the valves’ dynamics, the problem is modeled as a discounted Markov Decision Process (MDP) $\left(\mathcal{X},\mathcal{U},c,P,\gamma\right)$. Let $\Delta(\cdot)$ be the simplex on any given space ($\cdot$), and define $c:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ the cost function, $P:\mathcal{X}\times\mathcal{U}\to\Delta(\mathcal{X})$ the transition probabilities, and $\gamma\in\left[0,1\right)$ the discount factor. At each time step $t$, the agent receives an observation $x_{t}\in\mathcal{X}$ and selects a control $u_{t}\in\mathcal{U}$. The system is then updated with $x_{t+1}\sim P\left(\cdot|x_{t},u_{t}\right)$.
@@
The goal of this work is to develop a throttle valve controller $\pi:\mathcal{X}\to\Delta(\mathcal{U})$ that sets the angle of the valve’s plate to a chosen angle from any initial position. The controller must not only perform optimally in nominal operating conditions by minimizing cumulative costs but also demonstrate robustness under various real-world scenarios. Specifically, the controller must be able to reject external disturbances that may affect both the input and output signals, while exhibiting minimal overshoot. Bearing this information in mind, we associate the different components of the MDP with the valve. At time $t$, let $\alpha_{t}$ be the angle of the valve’s plate and $\alpha_{\text{ref}_{t}}$ be the (desired) reference angle, and:
@@
  • the controller must be optimal in all states, so the discount factor $\gamma$ is set to the high value of $0.99$;
@@ III-B Approximate Policy Iteration
This cycle of evaluation and improvement is commonly solved approximately with gradient descent. This framework, not requiring prior knowledge from the environment beyond interaction data, can be applied as such to any system as long as it possesses the Markov property. Nonetheless, the absence of initial information makes it difficult to collect meaningful data at the early stages of learning. This leads to poor estimates of the $Q$ function and the policy. As a result, agents necessitate extensive interactions to build accurate estimates and derive a near-optimal policy.
@@ IV Combining both methods
Figure 3: Comparison of agents under different scenarios with access to real outputs. Both figures show the evolution of angles over time with a reference change every $X$ seconds, set to 5 seconds for the left figures and 2.5 seconds for the right figures.
@@
Unlike the traditional RL policy $\pi_{\text{RL}}^{\theta}$ that provides a control input $u$, the perturbation $\xi^{\phi}_{\text{RL}}$ aims solely to enhance the PI controller’s performance. Thus, rather than defining the perturbation $\xi^{\phi}_{\text{RL}}$ over the entire control space $\mathcal{U}$, it is constrained to a smaller subset $S(\mathcal{U})$. This confines the search for control to a vicinity around the PI controller, where the optimal control is more likely to be found. In practice, to reduce the search space (originally $[0,0.6]$ or $[0,0.8]$) while keeping the policy differentiable, a multiplicative factor $\eta\in[0,1]$ is introduced that scales the control input by $\eta$.
The perturbation is learned similarly to a control input in traditional RL, that is by minimizing its associated $Q^{\omega}$ function. Note that the expectation in the Bellman operator is with respect to $\pi^{\phi_{k}}_{\text{PI-RL}}$, not $\pi_{\theta_{k}}$. This avoids the overestimating problem that may be encountered when the $Q$-function estimates observation-control pairs that are not described by the data set $\mathcal{D}$ [31, 26]. The process becomes the PI-Approximate Policy Iteration, where the APE-API steps are modified to be PI-APE and PI-API, defined as:
@@ V Comparison of the different controllers on the experimental test benches
Figure 4: Comparison of agents under noisy outputs or controls with varying standard deviations. Each bar represents the mean and standard deviation of the Mean-Squared-Error (MSE) between the reference and the angle across the $3$ valves.
We conduct a series of experiments on various tasks, including tracking a reference signal with sufficient time to adjust, adapting to quick reference changes, and demonstrating robustness with respect to noise in the control input or in the output observation.
@@ V-A Nominal comparison and reaction to quick reference changes
In all scenarios, we observe that the PI controller, while efficient, encounters difficulties in accurately tracking the reference signal. The controller requires time to accumulate errors before adjusting to the reference. Due to the system’s complexity and the dry friction effect, it may cause the valve’s plate to overshoot the target. The PI controller would then need additional time to adjust, and potentially select overly aggressive inputs once more. This results in oscillations around the reference angle, particularly evident in the first valve. The other valves also present such oscillations, especially for low-angle references. This reveals the need for more sophisticated controllers and not solely relying on a linear policy.
@@
On the other hand, both RL-based agents demonstrate remarkable success in precisely tracking the reference signal in all scenarios, including the most challenging ones. For example, they can track all reference values with a change every 2.5 seconds, which is twice as fast as the rate they were trained for. This level of performance is also observed for low-angle references that proved to be challenging for the PI controller. The controls provided by the RL-based agents are stable in all settings, with no instances of unstable behavior observed. In some cases, a slight tracking bias is observed where the agents reach a point close to but not precisely at the reference. This can be attributed to the RL agents’ goal of minimizing the $Q$ function, which is an expectation over the system’s stochastic dynamics. To address this, a more appropriate RL formulation may be needed, which considers higher-order moments of the cumulative future cost distribution or steady-state errors in a refined way. We leave this for future work. Despite these minor biases, both RL-based agents demonstrate a strong performance in all tasks for all valves.
@@ V-B Robustness to external noise
In comparison to the PI controller, both RL-based agents demonstrate superior performance over the PI controller in all scenarios. These RL-based agents have approximately the same performance in this setting, which is expected given both of them have been trained until convergence. It is also essential to note that, under specific conditions of noise, e.g., with an output or control noise of std $0.2$, the PI-RL agent exhibits a higher loss than the traditional RL agent. This result may be attributed to the fact that the PI-RL agent is centered around the PI controller which is inadequate in these scenarios: the reduced search space may not be sufficient to yield an optimal controller. Nevertheless, the perturbation $\xi_{RL}^{\phi}$ can improve the PI controller up to a certain point.
@@ V-C Sample efficiency
As previously shown, both the PI-RL and classical RL agents achieve similar performances in controlling the throttle valve. One important advantage of the combined PI and RL approach over the classical RL approach is that it requires fewer data to learn even though the PI controller may not be the optimal one. This can be seen in Figure 5, where we plot the cumulative costs over the training process for both controllers averaged over the $3$ different valves.
We observe that the PI-RL approach achieves a low cumulative cost faster than the RL approach, indicating that it was able to learn the control task more efficiently. This is because PI-RL leverages the prior knowledge of the PI controller to reduce the search space and guide the RL exploration, which helps to reduce the exploration time and improve the sample efficiency. After a sufficient amount of data, the traditional agent finally catches the performance of the combined agent.
@@
Figure 5: Learning curves of the different RL-based agents averaged on the $3$ valves. Agents are both trained with $4$ CPUs during $2500$ episodes, representing $3$ hours.
@@ VI Conclusion
In this work, we show that the combination of reinforcement learning and optimal control can be a powerful approach to designing near-optimal controllers for throttle valve systems. By adapting the recent area of Reinforcement Learning with Guides to this use case, we built a near-optimal controller that reduces the data requirement of traditional RL agents and overcomes the limitations of the PI controller. Our approach is validated on three valves with slightly different dynamics.
This provides evidence that the integration of Control Theory techniques with RL can lead to significant improvements in data efficiency, making it a promising avenue for future research in other complex systems.
ACKNOWLEDGMENT
The authors would like to thank Sophie CAPOBIANCO, Cheikh Saliou Fall NDIAYE and Adama DIABATE from Grenoble INP ENSE3 for their diligent efforts in realizing some of the experiments presented in this paper.